# Hidden State Poisoning Attacks against Mamba-based Language Models

### v1

@@ 4.1 Evaluated Models
We use our RoBench-25 to assess robustness of LLMs against adversarial prompt attacks (cf. HiSPA). We particularly focus on two models of the same parameter size and context window (2,048 tokens), trained using the same data (The Pile (gao2020pile)): one pure SSM model, Mamba-2.8b [^3] https://huggingface.co/state-spaces/mamba-2.8b-hf (gu_mamba_2024), and one pure Transformer, Pythia-2.8b [^4] https://huggingface.co/EleutherAI/pythia-2.8b (biderman_pythia_2023). Neither model has been finetuned. We refer to them as “mamba” and “pythia” hereafter. Additionally, we evaluate the full-precision Jamba-1.7-Mini [^5] https://huggingface.co/ai21labs/AI21-Jamba-Mini-1.7 (most recent release as of November 2025) (“jamba”, cf. §˜2) to test whether hybrid SSM-Transformer architectures inherit this vulnerability. As an instruction-tuned model, jamba serves as a strong baseline expected to achieve near-perfect scores in the absence of effective attacks. Experiments use a Tesla V100-SXM3-32GB for the small models and an NVIDIA H200 NVL for jamba.
@@ 4.2 Z-HiSPA Results
Table˜1 summarizes the results of RoBench-25 for mamba, pythia and jamba across 7 Z-HiSPA triggers (cf. App.˜B) and 4 configurations (with/without awareness and/or recovery instructions). Table˜1 clearly shows that HiSPA triggers have a dramatic impact on mamba’s performance (71% average CHSS drop from clean to HiSPA prompts with recovery instruction). In contrast, for Pythia-2.8b, the exact same configuration leads to an improvement (58% average CHSS increase). [^6] We hypothesize this is related to the effect that distractors may mitigate Transformers’ miscalibration in Question-Answering tasks (chhikara_mind_2025). This confirms both the severity of HiSPA attacks on Mamba, despite the lack of trigger optimization, and their specificity to its attention-free architecture. Jamba shows nearly no performance degradation under HiSPA attacks, as expected from its hybrid SSM-Transformer design.
@@ 4.3 M-HiSPA Results
While Z-HiSPAs demonstrate that zero-shot triggers can degrade Mamba, we now investigate whether optimized triggers can amplify this effect. As introduced in §˜3.3, M-HiSPAs assume white-box access, allowing us to optimize triggers directly against the model’s internals.
We employ a standard genetic algorithm (GA) (holland_adaptation_1992; goldberg_genetic_1988) with tournament selection, single-point crossover, and per-token mutation. The GA optimizes the loss function from eq.˜5 over the first block output embedding $\mathbf{o}(\mathbf{x})$, using 5 training and 5 held-out RoBench-25 samples. All runs surpassed the greedy beam search, driving the loss below -0.998 on this first-block objective. However, most GA-generated M-HiSPA triggers yield a smaller degradation on RoBench-25 than the Z-HiSPAs of §˜4.2. This mismatch between first-block loss minimization and downstream robustness collapse indicates that HiSPAs cannot be explained by early-block dynamics alone, as discussed further in §˜5.
One notable exception arises where the GA discovers a trigger containing the special token <|endoftext|>. This sequence pushes the first-block objective even further (loss -0.9998, Fig.˜3) and drives RoBench-25 performance for Mamba-2.8B down to the level of a random guesser. The same trigger also collapses Pythia-2.8B and Jamba-1.7-Mini to zero CHSS, despite the latter’s much larger scale, while all instruction-tuned Transformer baselines we tested (Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, SmolLM3-3B) resisted this attack. While HiSPAs are defined as attacks that exploit SSM recurrence, zhou_virtual_2024 show that special tokens can disrupt attention layers. Because this outlier trigger simultaneously leverages first block hidden-state saturation and the semantics of a special end-of-prompt marker, it likely reflects a broader, mixed failure mode that is not specific to recurrent hidden-state poisoning alone.
Jamba’s vulnerability to this hybrid attack already raises concerns about its suitability for deployment in adversarial environments. Here we focus on HiSPA-style triggers that do not rely on such special tokens. Future work will analyze these more powerful special-token triggers in detail, particularly in hybrid SSM-Transformer architectures, where SSM layers dominate.
@@ 5 Blockwise HiSPA Analysis
Block Pearson $r$ No Trigger Trigger A Trigger B Trigger C Trigger D Trigger E Trigger F Trigger G
@@
Table 2: Block number, correlation coefficient and corresponding $p$-value, and for each trigger (cf. App.˜B), the norm $\|\mathbf{o}^{(b)}_{t}\|_{2}$ for the 10 blocks most correlated with RoBench-25 performance.
@@
Figure 4: Relation between the block-29 output norm at the trigger and RoBench-25 scores, with linear fit, for different HiSPA triggers (A-G), cf. App.˜B.
@@ 5.1 Mechanistic Interpretation:
Our analysis from §˜3 and in this section indicates a two-stage pattern, comprising saturation of the hidden state at the first Mamba block and a second amplification in the middle blocks (28–37). This aligns with prior mechanistic studies of Mamba. sharma_locating_2024 localize factual associations in middle layers using causal tracing and rank-one editing, identifying an early-to-middle “enrichment” regime where subject representations are consolidated. ensign_investigating_2024 find that late layers act as retrieval bottlenecks, requiring well-structured representations from earlier blocks. rezaei_jafari_mambalrp_2024 show via MambaLRP that subtle SSM dynamics strongly influence which tokens are deemed relevant at output. Together, these findings suggest that blocks 28–37 constitute a critical transport corridor: HiSPA-induced norm amplification corrupts representations precisely where factual content is finalized, preventing downstream recovery even if late-layer circuits remain intact. The result is a two-stage failure mechanism (initial state saturation followed by mid-layer amplification) that systematically undermines information retention.
@@ 5.2 Implications for HiSPA Detection
Table 3: Attack success value (ASV) of PIAs on the updated Open-Prompt-Injections benchmark (liu_formalizing_2024), incl. absolute and relative differences from each model’s naive (no trigger) baseline ASV (smaller is better).
@@ 6 Connecting HiSPAs to PIAs
So far, we only discussed HiSPAs wherein no malicious instruction is explicitly provided to the model, and showed in §˜4 that the large, finetuned jamba model (Jamba-1.7-Mini, 52B) is nearly immune to Z-HiSPAs (zero-shot black-box HiSPAs) alone. However, the M-HiSPA optimized by our genetic metaheuristic completely prevents jamba from answering correctly while not impacting any of the evaluated finetuned Transformer models (from 3B to 70B, cf. §˜4.3), suggesting that the vulnerability is not limited to attention-free small models only.
M-HiSPAs tend to mimic some of the prompt injection attacks (PIAs) triggers studied in the Open-Prompt-Injections benchmark from liu_formalizing_2024, who demonstrated the effectiveness of simple PIAs against a wide range of LLMs, including GPT-4. After upgrading that benchmark to make it compatible with modern HuggingFace SSM accelerators, we evaluate jamba and two finetuned Llama models (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct, both full precision)777team_jamba-15_2024 compared jamba to 8B and 70B Llama models, we therefore use their most recent versions. against all PIA strategies from the benchmark, limiting ourselves to the 3 first datasets (sentiment analysis, spam detection and duplicate sentence reduction) for computational reasons. [^8] Since liu_formalizing_2024 found similar trends across all 7 datasets, we expect our conclusions to generalize.
Table˜3 lists the attack success value (ASV) for all tested configurations, as defined by liu_formalizing_2024. These range from directly injecting prompts without preceding triggers (“naive”), to combining the escape, ignore and fake completion strategies (“combine”). More details are provided in (liu_formalizing_2024).
In essence, “combine” adds a HiSPA-like trigger before the injected prompt, similar to our Z-HiSPA design. We therefore understand the relative difference between “naive” and “combine” as the impact of HiSPAs on PIAs, and find that Jamba’s ASV increases considerably from “naive” to “combine” (+25.3% absolute, $\times\,$1.516 relative). While the 8B Llama model experiences a milder degradation (+10.7% absolute, $\times\,$1.265 relative), the 70B Llama model rather gets less vulnerable when HiSPAs are added (-10.7% absolute, $\times\,$0.607 relative). This mirrors our findings in §˜4.2 (i.e., considerable degradation for mamba, improvement for pythia) at the LLM scale and with additional prompt injections. This observation further suggests that HiSPAs can have a significant impact on SSM models even in hybrid finetuned LLMs, motivating future work towards the development of dedicated defenses against them, and systematic evaluation of the robustness of SSM-based LLMs against HiSPAs and PIAs.
@@ 7 Conclusion and Further Work
State space models are rapidly transitioning from research prototypes to production components in long-context and efficiency-critical LLM deployments. We showed that the recurrent dynamics enabling their efficiency also constitute a security liability: Hidden State Poisoning Attacks (HiSPAs) exploit Mamba’s selectivity mechanism to irreversibly overwrite contextual information, causing severe performance degradation that purely attention-based models do not exhibit. Critically, this vulnerability does not disappear with scale or finetuning: it persists in the 52B-parameter Jamba-1.7-Mini and amplifies effectiveness of Prompt Injection Attacks compared to Transformer baselines.
@@
Beyond defensive detection techniques, this work opens several other research directions, including on the HiSPA vulnerability of different and emerging model families, on training-time regularization strategies to reduce susceptibility, or on extending RoBench-25 to multilingual and multi-hop reasoning settings, beyond the retrieval setting investigated in our work.
@@
8 Limitations
Architectural Coverage.
Our experiments focus on the original Mamba architecture and the Jamba hybrid family. Although (as stated in §˜2) Mamba is a very popular SSM design, other SSM architectures (like Mamba-2) have also gained traction recently. Their different mathematical formulations and parameterizations may affect their susceptibility to HiSPAs, and would merit dedicated analysis to cover the full spectrum of modern SSM-based LLMs. Extending our framework to these architectures is a natural avenue for future work.
Optimization of HiSPAs and computational resources.
Detection Not Implemented.
Additionally, §˜5.2 proposes norm-based HiSPA detection motivated by the strong block-29 correlation, but does not implement or evaluate such a detector. As discussed in §˜5, practical deployment would require threshold calibration, domain-specific tuning, and false-positive analysis, which we leave to future investigation.
@@ 9 Ethical Considerations
Appendix A Bounded step sizes and multi-step triggers
@@
Appendix B HiSPA Triggers Used
In this appendix, we list the different HiSPA triggers and defensive instructions used in our experiments (cf. §§˜3 and 4). Those triggers are not meant by any way to manipulate an LLM review of this paper.
B.1 HiSPA Triggers

### v3

@@ 4.1 Evaluated Models
We use our RoBench-25 to assess robustness of LLMs against adversarial prompt attacks (cf. HiSPA). We particularly focus on two models of the same parameter size and context window (2,048 tokens), trained using the same data (The Pile (Gao et al., 2020)): one pure SSM model, Mamba-2.8b [^6] https://huggingface.co/state-spaces/mamba-2.8b-hf (Gu and Dao, 2024), and one pure Transformer, Pythia-2.8b [^7] https://huggingface.co/EleutherAI/pythia-2.8b (Biderman et al., 2023). Neither model has been finetuned. We refer to them as “mamba” and “pythia” hereafter. Additionally, we evaluate the full-precision Jamba-1.7-Mini (“jamba”, cf. §˜2) to test whether hybrid SSM-Transformer architectures inherit this vulnerability. As an instruction-tuned model, jamba serves as a strong baseline expected to achieve near-perfect scores in the absence of effective attacks. Experiments use a Tesla V100-SXM3-32GB for the small models and an NVIDIA H200 NVL for jamba.
@@ 4.2 Z-HiSPA Results
Table˜1 summarizes the results of RoBench-25 for mamba, pythia and jamba across 7 Z-HiSPA triggers (cf. App.˜C) and 4 configurations (with/without awareness and/or recovery instructions). Table˜1 clearly shows that HiSPA triggers have a dramatic impact on mamba’s performance (71% average CHSS drop from clean to HiSPA prompts with recovery instruction). In contrast, for Pythia-2.8b, the exact same configuration leads to an improvement (58% average CHSS increase). [^8] We hypothesize this is related to the effect that distractors may mitigate Transformers’ miscalibration in Question-Answering tasks (Chhikara, 2025). While Mamba’s clean CHSS is modest, reflecting base models’ inherent difficulty at structured retrieval (cf. App.˜A), the differential pattern (71% degradation for Mamba vs. 58% improvement for Pythia under identical triggers) confirms the architecture-specificity of HiSPAs. We show in App.˜H that similar results hold for Mamba-2 and for the Nemotron-3-Nano hybrid. Jamba, which achieves near-perfect clean CHSS, shows no Z-HiSPA degradation, confirming its hybrid design’s resilience.
@@ 4.3 M-HiSPA Results
We now investigate whether optimized M-HiSPA triggers (cf. §˜3.3) can amplify the observed Z-HiSPA degradation.
We employ a standard genetic algorithm (GA) (Holland, 1992; Grefenstette, 1993) with tournament selection, single-point crossover, and per-token mutation. The GA optimizes the loss function from eq.˜5, using 5 training and 5 held-out RoBench-25 samples. All runs surpassed the greedy beam search, driving the loss below -0.998 on this first-block objective. However, most GA-generated M-HiSPA triggers yield smaller RoBench-25 degradation than Z-HiSPAs, despite achieving lower first-block loss. This indicates that first-block saturation is necessary but not sufficient for effective HiSPAs; §˜5 shows that mid-layer norm amplification provides the missing explanatory link.
One notable exception arises where the GA discovers a trigger containing the special token <|endoftext|>. This sequence pushes the first-block objective even further (loss -0.9998, Fig.˜4) and collapses Mamba-2.8B to random-guessing level on RoBench-25. The same trigger also collapses Pythia-2.8B and Jamba-1.7-Mini to zero CHSS, despite the latter’s much larger scale, while all instruction-tuned Transformer baselines we tested (Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, SmolLM3-3B) resisted this attack. While HiSPAs are defined as attacks that exploit SSM recurrence, Zhou et al. (2024) show that special tokens can disrupt attention layers. Because this outlier trigger simultaneously leverages first block hidden-state saturation and the semantics of a special end-of-prompt marker, it likely reflects a broader, mixed failure mode that is not specific to recurrent hidden-state poisoning alone.
The SSM-specific vulnerability of Jamba at scale is instead demonstrated in §˜6, where non-special-token triggers amplify PIA success on Jamba ($\times$ 1.516) while reducing it on Transformer baselines. We hereafter focus on HiSPA-style triggers without special tokens. Future work will analyze special-token triggers in detail, particularly in hybrid SSM-Transformer architectures.
@@ 5 Blockwise HiSPA Analysis
Block Pearson $r$ No Trigger Tri. A Tri. B Tri. C Tri. D Tri. E Tri. F Tri. G
@@
Table 2: Block number, correlation coefficient and corresponding $p$-value, and for each trigger (cf. App.˜C), the norm $\|\mathbf{o}^{(b)}_{t}\|_{2}$ for the 10 blocks most correlated with RoBench-25 performance.
@@
Figure 4: Relation between the block-29 output norm at the trigger and RoBench-25 scores, with linear fit, for different HiSPA triggers (A-G), cf. App.˜C.
@@ 5.1 Mechanistic Interpretation:
Our analysis from §˜3 and in this section indicates a two-stage pattern, comprising saturation of the hidden state at the first Mamba block and a second amplification in the middle blocks (28–37). This aligns with prior mechanistic studies of Mamba. Sharma et al. (2024) localize factual associations in middle layers using causal tracing and rank-one editing, identifying an early-to-middle “enrichment” regime where subject representations are consolidated. Ensign and Garriga-Alonso (2024) find that late layers act as retrieval bottlenecks, requiring well-structured representations from earlier blocks. Rezaei Jafari et al. (2024) show via MambaLRP that subtle SSM dynamics strongly influence which tokens are deemed relevant at output. Together, these findings suggest that blocks 28–37 constitute a critical transport corridor: HiSPA-induced norm amplification corrupts representations precisely where factual content is finalized, preventing downstream recovery even if late-layer circuits remain intact. The result is a two-stage failure mechanism (initial state saturation followed by mid-layer amplification) that systematically undermines information retention. Whether a similar two-stage pattern occurs in Mamba-2 is investigated in App.˜H.
@@ 5.2 Implications for HiSPA Detection
Table 3: Attack success value (ASV) of PIAs on the updated Open-Prompt-Injections benchmark (Liu et al., 2024), incl. absolute and relative differences from each model’s naive (no trigger) baseline ASV (smaller is better).
@@ 6 Connecting HiSPAs to PIAs
So far, we discussed HiSPAs without explicit malicious payloads (e.g., in the context of paper injection, demanding a positive review), showing (§˜4) that jamba is nearly immune to Z-HiSPAs alone but collapses under the M-HiSPA trigger that does not affect finetuned Transformers (3B–70B, cf. §˜4.3), suggesting the vulnerability extends beyond small attention-free models.
HiSPAs tend to mimic PIA triggers from the Open-Prompt-Injections benchmark (Liu et al., 2024), which demonstrated the effectiveness of simple PIAs against a wide range of LLMs, including GPT-4. After upgrading that benchmark to make it compatible with modern HuggingFace SSM accelerators, we evaluate jamba and two finetuned Llama models (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct, both full precision) [^9] Team et al. (2024) compared jamba to 8B and 70B Llama models, we therefore use their most recent versions. against all PIA strategies from the benchmark, limiting ourselves to the 3 first datasets (sentiment analysis, spam detection and duplicate sentence reduction) for computational reasons. [^10] Since Liu et al. (2024) found similar trends across all 7 datasets, we expect our conclusions to generalize.
Table˜3 lists the attack success value (ASV) for all tested configurations, as defined by Liu et al. (2024). These range from directly injecting prompts without preceding triggers (“naive”), to combining the escape, ignore and fake completion strategies (“combine”). More details are provided in (Liu et al., 2024).
In essence, “combine” adds a HiSPA-like trigger before the injected prompt, similar to our Z-HiSPA design. We therefore understand the relative difference between “naive” and “combine” as the impact of HiSPAs on PIAs, and find that Jamba’s ASV increases considerably from “naive” to “combine” (+25.3% absolute, $\times\,$1.516 relative). While the 8B Llama model experiences a milder degradation (+10.7% absolute, $\times\,$1.265 relative), the 70B Llama model rather gets less vulnerable when HiSPAs are added (-10.7% absolute, $\times\,$0.607 relative). This mirrors our findings in §˜4.2 (i.e., considerable degradation for mamba, improvement for pythia) at the LLM scale, and, crucially, provides clean, SSM-specific evidence for jamba’s vulnerability using non-special-token triggers, complementing the RoBench-25 results where the only effective jamba trigger relies on the <|endoftext|> special token. This confirms that HiSPAs have a significant impact on SSM models even in hybrid finetuned LLMs, providing cross-task validation of HiSPA vulnerability beyond RoBench-25’s retrieval setting, and motivating dedicated defenses and systematic robustness evaluation of SSM-based LLMs.
@@ 7 Conclusion and Further Work
State space models are rapidly transitioning from research prototypes to production components in long-context and efficiency-critical LLM deployments. We showed that the recurrent dynamics enabling their efficiency also constitute a security liability: Hidden State Poisoning Attacks (HiSPAs) exploit the selectivity mechanism of Mamba and Mamba-2 to irreversibly overwrite contextual information, causing severe performance degradation that purely attention-based models do not exhibit. Critically, this vulnerability does not disappear with scale or finetuning: it persists in the 52B-parameter Jamba-1.7-Mini and amplifies effectiveness of Prompt Injection Attacks compared to Transformer baselines.
@@
Beyond defensive detection techniques, this work opens several other research directions, including on the HiSPA vulnerability of further emerging model families beyond the Mamba and Mamba-2 architectures studied here, on training-time regularization strategies to reduce susceptibility, or on extending RoBench-25 to multilingual and multi-hop reasoning settings, beyond the retrieval setting investigated in our work.
@@ 8 Ethical Considerations
Acknowledgements
We disclose the use of generative artificial intelligence tools during the preparation of this work, in accordance with the Responsible NLP Checklist. Apart from language models being the topic of study of this work, AI-based systems were also used to assist with (i) writing and debugging code, (ii) exploratory literature search and summarization, and (iii) rephrasing and editing of selected passages of the manuscript for clarity and conciseness. In addition, code development was assisted by large language models including ChatGPT, Grok, and Claude.
All AI-assisted outputs were reviewed, validated, and, where necessary, modified by the authors. All scientific decisions, experimental design, data analysis, interpretation of results, and conclusions were performed entirely by the authors, who take full responsibility for the content of the manuscript and for any errors or omissions.
Appendix A Limitations
Baseline RoBench-25 scores.
Although the baseline CHSS values for some models may appear low, each reported score is averaged over 120 distinct prompt configurations (240 queries), which substantially reduces the likelihood that the observed differences arise from random noise. As a result, the relative performance gaps across experimental conditions, especially in cases where CHSS approaches zero, remain meaningful and are unlikely to be sampling artifacts.
Several considerations explain why these baselines do not undermine our central claims. First, we deliberately chose non-instruction-tuned base models (Mamba-2.8b, Pythia-2.8b) to isolate the architectural effect of SSM recurrence from confounding factors such as instruction-following training, RLHF, or safety fine-tuning. Base models are inherently weaker at structured question answering, which explains the moderate clean CHSS. Second, the key evidence for HiSPA vulnerability is not Mamba’s absolute CHSS but the differential comparison: identical triggers that reduce Mamba’s CHSS by 71% simultaneously increase Pythia’s by 58%, a divergence that cannot be attributed to weak baselines. Third, we validate HiSPA effects on models with strong clean baselines: Jamba-1.7-Mini (clean CHSS $\approx$ 1.0) collapses under the M-HiSPA trigger (§˜4.3), and HiSPA-like prefixes substantially amplify PIA success rates on Jamba compared to Transformer baselines (§˜6). Fourth, the Open-Prompt-Injections evaluation (§˜6) provides a second, independent task confirming the SSM-specific amplification effect on a production-scale instruction-tuned model.
We nevertheless acknowledge that a benchmark yielding higher clean CHSS for Mamba and Mamba-2 would provide a more convincing demonstration of the phenomenon. Despite this limitation, the consistency of the degradations across triggers, configurations, and random seeds provides strong empirical evidence for the HiSPA effect under the evaluated setup, and confirms our theoretical demonstrations.
Optimization of HiSPAs and computational resources.
Detection Not Implemented.
Additionally, §˜5.2 proposes norm-based HiSPA detection motivated by the strong block-29 correlation, but does not implement or evaluate such a detector. As discussed in §˜5, practical deployment would require threshold calibration, domain-specific tuning, and false-positive analysis, which we leave to future investigation. We note, however, that Table˜2 provides suggestive evidence for feasibility: block-29 norms for triggered prompts (range 9.80–12.06) consistently exceed the clean-prompt norm (9.49), with the gap widening for more effective triggers (e.g., trigger G at 12.06 vs. 9.49 clean). A meaningful separation between clean and triggered norm distributions is a prerequisite for threshold-based detection, and the observed gap is encouraging.
Optimization–performance mismatch.
As discussed in §§˜4.3 and 5, GA-optimized M-HiSPAs achieve lower first-block loss than Z-HiSPAs yet often produce smaller RoBench-25 degradation. Our blockwise analysis (§˜5) provides a partial explanation: effective HiSPAs require not only first-block state saturation but also norm amplification in the mid-layer band (blocks 28–37). Optimizing exclusively on first-block representations may find triggers that saturate the initial state without inducing the necessary downstream amplification. A multi-objective optimization targeting both first-block saturation and mid-layer norm amplification could potentially yield stronger M-HiSPAs, but would substantially increase computational requirements. We note that this mismatch, rather than weakening our theoretical framework, actually strengthens it by revealing the two-stage nature of the poisoning mechanism: it shows that hidden-state corruption propagates through a specific computational pathway that must be fully engaged for the attack to succeed.
Special-token confound in Jamba evaluation.
The M-HiSPA trigger that collapses Jamba-1.7-Mini on RoBench-25 contains the <|endoftext|> special token, which also collapses Pythia-2.8b (a pure Transformer), preventing clean attribution to SSM-specific dynamics. The SSM-specific evidence for Jamba’s vulnerability comes instead from the Open-Prompt-Injections evaluation (§˜6), where non-special-token HiSPA-like triggers amplify attack success on Jamba ($\times$ 1.516) while reducing it on Llama-3.3-70B. These two lines of evidence are complementary: RoBench-25 demonstrates that Jamba can be collapsed (even if through a mixed mechanism), while the PIA evaluation demonstrates a clean, SSM-specific amplification effect at production scale. Developing effective M-HiSPA triggers for Jamba that do not rely on special tokens remains an open challenge.
Appendix B Bounded step sizes and multi-step triggers
@@
Remark.
The derivation above requires only submultiplicativity of the norm and a uniform contraction bound $\rho<1$; it does not depend on $\bar{\mathbf{A}}_{t}$ being a general diagonal. In particular, when $\bar{\mathbf{A}}_{t}=\alpha_{t}I_{N}$ (scalar-identity, as in Mamba-2), all bounds carry through verbatim. The full adaptation to the Mamba-2 parameterization is provided in App.˜H.
Appendix C HiSPA Triggers Used
In this appendix, we list the different HiSPA triggers and defensive instructions used in our experiments (cf. §§˜3 and 4). Those triggers are not meant in any way to manipulate an LLM review of this paper.
C.1 HiSPA Triggers